Can anyone explain to me why I've seen the equation

Bias= 2^exponentbits -1 and

Bias= (2^exponentbits -1)-1

I'm confused on what equation to use. Can you please clarify? Thanks
Answer #1

Ans:

Bias = ( 2exponent_bits - 1)-1

It is the correct expression to find Bias

Because ,

Let's take single precision (32-bit), among 32 bits we have 1 sign bit, 8 exponent bits and 23 magnitude bits.

So, the Bias =(28-1)-1 = 27 -1 =128 -1 = 127

Actually the Bias for single precision is 127 only.

So, the correct equation is

Bias = (2exponent_bits - 1) - 1